A common mistake when solving equations is the following:
The equation: 2(x-2) = x + 3
First step in solving: 2x- 2= x + 3
How do I write a explanation of what error has been made. What could be done to
avoid this error?

Answer by tutorcecilia(2152)


The mistake is that all of the terms within the parenthesis should have been multiplied by the "2" from the outside of the parenthesis.
Corrected equation:
2(x-2)=x+3
(2)x-2(2)=x+3 [distribute or multipliy all of the terms within the parenthesis by "2"]
.
2x-4=x+3